# Quantum anomalous, quantum spin, and quantum valley Hall effects in pentalayer rhombohedral graphene

### v1

@@ Abstract
Recent experiments in pentalayer rhombohedral graphene moiré superlattice have observed the quantum anomalous Hall effect at the moiré filling factor $\nu=1$ and various fractions. These effects result from a flat Chern band induced by electron-electron interactions. In this Letter, we show that the many-body effects at $\nu=2$ give rise to the quantum spin Hall and quantum valley Hall states, even without spin-orbit couplings or valley-dependent potentials, in addition to the quantum anomalous Hall state. These three topological states can be selectively induced by applying and tilting a magnetic field. Furthermore, we demonstrate that at $\nu=3$ and $4$, a combination of the Wigner-like crystal state and topological states can be the ground state. Consequently, the relation between states at the filling factor $\nu$ and its particle-hole counterpart $4-\nu$ breaks down, which contrasts with the conventional quantum Hall effect in graphene.
Introduction. —
@@
The recent observation of the integer and fractional quantum anomalous Hall (QAH) effects in moiré materials [1, 2, 3, 4, 5] is a significant advancement in condensed matter physics. Moiré materials are a platform of strong interaction effects [6, 7], resulting in a variety of symmetry-broken phases such as superconductivity and Wigner crystals [8, 9, 10, 11, 12], as well as topological phases [13, 14, 15, 16, 17, 18]. The QAH effect arises from the interplay of symmetry breaking and topology, where the Hall conductance is quantized even without external magnetic fields [19, 20, 21, 22]. Among moiré materials, the integer and fractional QAH effects has been first observed in twisted bilayer MoTe2 [1, 2, 3, 4]. Recently, the QAH states in pentalayer rhombohedral graphene on hexagonal boron nitride (hBN) at the moiré filling factor $\nu=1$ and some fractions have been reported [5]. Following up this breakthrough, intense theoretical effort has been made  [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45].
@@
Chern bands in pentalayer rhombohedral graphene on hBN emerge from the interplay of three factors: layer stacking, moiré structure, and many-body interactions. In turn, the band touching at the charge-neutral point of multilayer rhombohedral graphene [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56] becomes flatter ($\sim k^{N_{L}}$ with the wave number $k$) as the number of stacked layer $N_{L}$ increases [57]. Moiré superlattice structure due to lattice mismatch between pentalayer graphene and hBN [58, 59] reduces the graphene’s Brillouin zone into the moiré one, making the band width smaller. Finally, the electron-electron interactions single out a Chern band from the rest of the moiré bands. Specifically, the Hartree-Fock (HF) analysis revealed the emergence of an isolated and nearly flat Chern band at $\nu=1$ [31, 32, 33, 34], which is absent without interaction. This mechanism is supposed to underlie the observed $\nu=1$ and fractional QAH states [5, 31, 32, 33].
@@
In general, quantum Hall physics [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72] is based on noninteracting topological bands like the Landau level as a starting point, including integer and fractional Chern insulators [73, 74, 75, 76, 77, 18]. However, pentalayer rhombohedral graphene on hBN falls outside of this framework, suggesting potentials for interaction-driven exotic phases beyond the conventional quantum Hall paradigm even at integer fillings. This motivates us to explore pentalayer rhombohedral graphene on hBN at integer fillings other than $\nu=1$, which have not yet been reported in experiments under a sufficiently strong displacement field [5].
@@
In this Letter, we show that pentalayer rhombohedral graphene at $\nu=2$ can yield the quantum spin Hall (QSH) and quantum valley Hall (QVH) states, even without spin-orbit couplings or valley-dependent potentials, in addition to the QAH state. They emerge purely from many-body effects, as in the topological Mott insulator [78]. Our self-consistent HF calculation shows that the QAH, QSH, and QVH states at $\nu=2$ are degenerate, while they are distinguishable by the magnitude of magnetization. This difference allows selective induction of one of the three states by applying and tilting magnetic fields, as shown in Fig. 1(a). The appearance of the three states can be understood from the analogy of $\nu=1$: the emergent Chern band at $\nu=1$ spontaneously selects the valley $K,K^{\prime}$ and spin $\uparrow$,$\downarrow$ , ↓, where the Chern number is $+1$ ($-1$) for $K$ ($K^{\prime}$[31, 32, 33, 34]. The ground state at $\nu=2$ is composed of two Chern bands out of $(K\!\!\uparrow,K\!\!\downarrow,K^{\prime}\!\!\uparrow,K^{\prime}\!\!\downarrow)$ [see Fig. 1(b)]. Different combinations carry different topological numbers, leading to the QAH, QSH, and QVH states. All possible cases are summarized in Table 1.
@@
Figure 1: (a) Lowest energy state as a function of in-plane and out-of-plane components $(B_{\perp},B_{\parallel})$ of the magnetic field ${\bm{B}}$. Boundaries (I)-(III) are identified as follows: (I) QAH vs QSH: $B_{\perp}=B_{C}$. The critical magnetic field $B_{C}$ is defined so that $B_{C}\times M_{\text{orb}}^{z\text{(QAH)}}$, where $M_{\text{orb}}^{z}$ is defined in Eq. (1), is equivalent to the the magnetostatic energy of the QAH state. (II) QVH vs QSH: $|{\bm{B}}|=\alpha B_{C}$ with $\alpha=4.9$ [see Eq. (2)]. (III) QVH vs QAH: $\alpha B_{\perp}=|{\bm{B}}|$. (b) Schematic of the band structure at $\nu=2$. The HF interaction induces two isolated bands. As numerically demonstrated below, the selection of the valley and spin – specifically, any two from $(K\!\!\uparrow,K\!\!\downarrow,K^{\prime}\!\!\uparrow,K^{\prime}\!\!\downarrow)$ – does not affect the total energy. All possible occupancies are summarized in Table 1.
@@
Remarkably, the empirical rule of forming the lowest energy state by occupying the Chern bands breaks down at $\nu=3$ and 4. We find that at these fillings, a combination of the Wigner-like crystal (WLC) state and topological states can be the ground state. This happens because their spatial distributions of the charge density interlock with each other, that is, WLC’s localized density fits into low-density spots of topological states. Consequently, the ground states at $\nu=3$ and $4$ are not necessarily the particle-hole counterparts of $14-\nu=$ 1 and 00 within the Chern band subspace consisting of $(K\!\!\uparrow,K\!\!\downarrow,K^{\prime}\!\!\uparrow,K^{\prime}\!\!\downarrow)$. This property is a many-body effect as well, and contrasts with the conventional quantum Hall effect in graphene.
Model —
@@
For numerical analysis, we use an effective continuum model [31, 33]. First we consider the tight-binding Hamiltonian of pentalayer rhombohedral graphene, including both intralayer and interlayer hoppings and the interlayer potential difference [See Sec. S1 in Supplementary Materials (SM) [79] for more details]. Employing a standard approach with a continuum limit, we obtain an effective Hamiltonian near the charge neutral point, $H_{R5G}=\sum_{{\bm{k}}}\bm{c}^{\dagger}({\bm{k}})h_{R5G}({\bm{k}})\bm{c}({\bm{k}})$ for a given valley and spin, where $\bm{c}^{\dagger}=(c_{A_{1}}^{\dagger},c_{B_{1}}^{\dagger},c_{A_{2}}^{\dagger},\ldots,c_{B_{5}}^{\dagger})$ and $c^{\dagger}_{X_{l}}$ is a creation operator for the sublattice $X=A,B$ on the layer $l=1,2,\ldots,5$. Here, $h_{R5G}({\bm{k}})$ is a 10-dimensional matrix. System parameters are set to match experimental conditions for the $\nu=1$ QAH effect [80, 31, 33, 5].
@@
Table 1: Possible occupancy of the four degrees of freedom at $\nu=2$, yielding ${}_{4}C_{2}=6$ states. The charge Chern number $C_{c}$ sums the band Chern numbers, while the spin (valley) Chern number $C_{s}$ $(C_{v})$ calculates the difference regarding the spin (valley). [Here, the band Chern number for the valley $K$ ($K^{\prime}$) is $+1$ ($-1$).] There are three types of states: quantum anomalous Hall (QAH), quantum spin Hall (QSH), and quantum valley Hall (QVH) states. The result for magnetization in Fig. 3 is also summarized. ($\circ$ indicates finite magnetization.)
@@ 6 $\times$ $\checkmark$ $\times$ $\checkmark$ (0,0,2)
Stacking graphene on hBN induces a moiré superlattice structure due to lattice mismatch. Here, we set the twist angle to $0.77^{\circ}$ to align with the experiment in Ref. 5. Within the effective continuum model, the effect of hBN is represented by a local potential $v({\bm{r}})$ within the bottom graphene subspace [58, 59] (see Sec. S2 in SM [79] for more details). Its second quantized form is $V_{\text{hBN}}=\sum_{{\bm{k}}}\sum_{m_{1}m_{2}}\tilde{\bm{c}}^{\dagger}({\bm{k}}+m_{1}\bm{G}_{1}+m_{2}\bm{G}_{2})v(m_{1},m_{2})\tilde{\bm{c}}({\bm{k}})$, where $\tilde{\bm{c}}^{\dagger}=(c_{A_{1}}^{\dagger},c_{B_{1}}^{\dagger})$, $\bm{G}_{1}$ and $\bm{G}_{2}$ the moiré reciprocal lattice vectors, and $v(m_{1},m_{2})$ is the Fourier coefficient of $v({\bm{r}})$. In the numerical calculations, the summation $\sum_{{\bm{k}}}\sum_{m_{1}m_{2}}$ is confined to the first and second moiré Brillouin zones.
@@
Now we obtained an effective noninteracting Hamiltonian as $H_{R5G}+V_{\text{hBN}}$. Then we add the HF interaction to incorporate many-body effects. The original interaction term is $H_{\text{int}}=(1/2S)\sum_{{\bm{k}}{\bm{k}}^{\prime}{\bm{q}}}\sum_{ZZ^{\prime}}V_{C}({\bm{q}})c^{\dagger}_{Z}({\bm{k}}+{\bm{q}})c^{\dagger}_{Z^{\prime}}({\bm{k}}^{\prime}-{\bm{q}})c_{Z^{\prime}}({\bm{k}}^{\prime})c_{Z}({\bm{k}})$, where $S$ is the are of the system, and $Z$ and $Z^{\prime}$ represent the spin, valley, sublattice, and layer indices. Here, we employ a dual gate-screened Coulomb interaction $V_{C}({\bm{q}})=e^{2}/(2\epsilon_{0}\epsilon_{r}q)\tanh(qd_{s})$ for ${\bm{q}}\neq{\bm{0}}$ and $V_{C}({\bm{q}}={\bm{0}})=0$, with $d_{s}=25$nm as the gate distance and $\epsilon_{r}=5$ as the dielectric constant [33]. See Sec. S3 in SM [79] for more details of the HF interaction. To reduce the computational cost, the following calculations are performed within the subspace of the first three noninteracting conduction bands for each valley and spin. We consider $24\times 24$ moiré unit cells unless otherwise stated.
Band structure —
@@
Our self-consistent HF calculation assumes the spin and valley conservation. This assumption regarding spin is valid because of the SU(2) spin-rotational symmetry of the HF Hamiltonian. Similarly, the assumption regarding the valley is justified since superimposing $K$ and $K^{\prime}$ gains additional energy [34].
@@
Figure 2: Band structures at $\nu=2$ of the lowest energy HF solutions. The Greek letters represent the high-symmetric points in the moiré Brillouin zone. Both figures provide the same many-body energy. The colors and line types indicate the valley and spin. (a) and (b) correspond to the states #3(QSH) and #1(QAH) in Table 1. In the HF calculation in (b), we introduce a small “valley Zeeman term” to favor the valley $K$ (see footnote [81]).
@@
Figures 2(a) and (b) show the HF band structures of the degenerate lowest energy solutions calculated by fixing the particle number with $\nu=2$. In Fig. 2(a) [(b)], the occupied two lowest bands, distinctly separated from other conduction bands, exhibit $(K\!\!\uparrow,K^{\prime}\!\!\downarrow)$ [$(K\!\!\uparrow,K\!\!\downarrow)$] with the band Chern numbers $C=(1,-1)$ [$C=(1,1)$], corresponding to the state #3(QSH) [#1(QAH)] in Table 1. Because our HF Hamiltonian maintains symmetry for spin flipping for each valley, these two results imply that the QAH, QSH, and QVH states in Table 1 are all the degenerate ground states. (For example, the state #5(QVH) is obtained by flipping the spin for $K^{\prime}$ in the state #3.)
Lowest energy state —

### v2

# Quantum Anomalous, Spin, and Valley Hall Effects in Pentalayer Rhombohedral Graphene Moiré Superlattices

@@ Abstract
Recent experiments on pentalayer rhombohedral graphene moiré superlattices have observed the quantum anomalous Hall effect at moiré filling factor of $\nu=1$ and various fractional values. These phenomena are attributed to a flat Chern band induced by electron-electron interactions. In this study, we demonstrate that at $\nu=2$, many-body effects can lead to the emergence of quantum spin Hall and quantum valley Hall states, in addition to the quantum anomalous Hall state, even in the absence of spin-orbit coupling or valley-dependent potentials. These three topological states can be selectively induced by the application and manipulation of a magnetic field. Furthermore, we show that at $\nu=3$ and $4$, the ground state can be a combination of topologically trivial and nontrivial states, unlike the cases of $\nu=1$ and 2. This contrasts with the conventional quantum Hall effect in graphene where the ground state at filling factor $\nu$ is given as the particle-hole counterpart at $4-\nu$.
I Introduction
@@
The recent discovery of both integer and fractional quantum anomalous Hall (QAH) effects in moiré materials represents a major milestone in condensed matter physics [1, 2, 3, 4, 5]. Moiré materials provide a fertile ground for exploring strong interaction effects [6, 7], leading to the emergence of various symmetry-broken phases such as superconductivity and Wigner crystals [8, 9, 10, 11, 12], as well as topological phases [13, 14, 15, 16, 17, 18]. The QAH effect is particularly intriguing because it arises from the interplay between symmetry breaking and topology, resulting in a quantized Hall conductance even in the absence of external magnetic fields [19, 20, 21, 22]. Both integer and fractional QAH effects were first observed in twisted bilayer MoTe2 [1, 2, 3, 4], marking a significant achievement in the study of moiré materials. More recently, these phenomena have been reported in pentalayer rhombohedral graphene on hexagonal boron nitride (hBN) at a moiré filling factor of $\nu=1$ and at various fractional values [5]. This breakthrough has spurred extensive theoretical investigations [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45].
@@
The emergence of Chern bands in pentalayer rhombohedral graphene on hBN can be attributed to the combined effects of layer stacking, moiré structure, and many-body interactions [31, 32, 33, 34, 35]. The band touching at the charge-neutral point of multilayer rhombohedral graphene  [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58] becomes increasingly flat (approximately $\sim k^{N_{L}}$ with wave number $k$) as the number of stacked layers $N_{L}$ increases [59]. The moiré superlattice structure, resulting from the lattice mismatch between pentalayer graphene and hBN [60, 61], reduces the graphene’s Brillouin zone (BZ) into moiré BZ, narrowing the band width. Ultimately, electron-electron interactions play a crucial role in isolating a Chern band from the rest of the moiré bands. Specifically, Hartree-Fock (HF) analysis has revealed the emergence of an isolated and nearly flat Chern band at $\nu=1$ [31, 32, 33, 34, 35], which does not appear in the absence of electron interactions. This mechanism is believed to underlie the observed $\nu=1$ and fractional QAH states [5, 31, 32, 33, 34, 35].
@@
Traditionally, quantum Hall physics [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 18] relies on non-interacting topological bands, such as Landau levels, as a foundation. However, pentalayer rhombohedral graphene on hBN deviates from this noninteracting framework, indicating the potential for interaction-driven exotic phases beyond the conventional quantum Hall paradigm, even at integer fillings. This observation prompts an investigation into pentalayer rhombohedral graphene on hBN at integer fillings other than $\nu=1$, which have not yet been explored experimentally under a sufficiently strong displacement field [5].
@@
In this study, we demonstrate that pentalayer rhombohedral graphene at $\nu=2$ can give rise to quantum spin Hall (QSH) and quantum valley Hall (QVH) states, in addition to the QAH state, even in the absence of spin-orbit coupling or valley-dependent potentials. These states emerge purely from many-body effects, which is a kind of a topological Mott insulator proposed in Ref. [80]. Our self-consistent HF calculations show that the QAH, QSH, and QVH states at $\nu=2$ are degenerate but distinguishable by their magnetization magnitudes. This distinction allows for the selective induction of one of these states by applying and tilting magnetic fields, as illustrated in Fig. 1(a). The emergence of these three states can be understood by analogy to the $\nu=1$ case: the emergent Chern band at $\nu=1$ spontaneously selects the valley $K$ or $K^{\prime}$ and spin $\uparrow$,$\downarrow$ , ↓, where the band Chern number is +1 for $K$ and -1 for $K^{\prime}$ [31, 32, 33, 34, 35]. At $\nu=2$, the ground state comprises two Chern bands selected from $(K\!\!\uparrow,K\!\!\downarrow,K^{\prime}\!\!\uparrow,K^{\prime}\!\!\downarrow)$ [see Fig. 1(b)]. Different combinations of the four degrees of freedom result in distinct topological states, leading to QAH, QSH, and QVH effects. The possible combinations are detailed in Table 1.
@@
Figure 1: (a) Lowest energy state as a function of the in-plane and out-of-plane components $(B_{\parallel},B_{\perp})$ of the magnetic field ${\bm{B}}$. The boundaries are defined as follows: (I) QAH versus QSH transition occurs at $B_{\perp}=B_{C}$, where $B_{C}$ is the critical magnetic field such that $B_{C}\times M_{\text{orb}}^{z\text{(QAH)}}$, with $M_{\text{orb}}^{z}$ defined in Eq. (1), equals the magnetostatic energy of the QAH state. (II) QVH versus QSH boundary at $|{\bm{B}}|=\alpha B_{C}$ with $\alpha=4.9$ [refer to Eq. (2)]. (III) QVH versus QAH transition described by $\alpha B_{\perp}=|{\bm{B}}|$. (b) Schematic of the band structure with and without the Hartree-Fock (HF) interaction at $\nu=2$. The HF interaction results in two isolated bands, as demonstrated in Fig. 2 below. The choice among the valley and spin – $(K\!\!\uparrow,K\!\!\downarrow,K^{\prime}\!\!\uparrow,K^{\prime}\!\!\downarrow)$ – does not influence the total energy. All possible occupancies are summarized in Table 1.
@@
Notably, the empirical rule of forming the lowest energy state by occupying the Chern bands does not hold at $\nu=3$ and $\nu=4$. At these filling factors, the ground state can be a combination of topologically trivial and nontrivial states. This occurs because the spatial charge density distributions of these states interlock; the localized density of the trivial state fits into the low-density regions of the topological states. This phenomenon, driven by many-body interactions, contrasts with the conventional quantum Hall effect in graphene where the ground state at filling factor $\nu$ is given as the particle-hole counterpart at $4-\nu$.
II Model
@@
For numerical analysis, an effective continuum model is used [31, 33]. The starting point is the tight-binding Hamiltonian of pentalayer rhombohedral graphene, including both intralayer and interlayer hoppings as well as the interlayer potential difference [see Appendix A]. Using a standard approach to reach the continuum limit, the effective Hamiltonian near the charge-neutral point is derived: $H_{R5G}=\sum_{{\bm{k}}}\bm{c}^{\dagger}({\bm{k}})h_{R5G}({\bm{k}})\bm{c}({\bm{k}})$ for a given valley and spin, where $\bm{c}^{\dagger}=(c_{A_{1}}^{\dagger},c_{B_{1}}^{\dagger},c_{A_{2}}^{\dagger},\ldots,c_{B_{5}}^{\dagger})$ and $c^{\dagger}_{X_{l}}$ is a creation operator for the sublattice $X=A,B$ on the layer $l=1,2,\ldots,5$. The matrix $h_{R5G}({\bm{k}})$ is ten-dimensional. System parameters are set to match experimental conditions for the $\nu=1$ QAH effect [81, 31, 33, 5]; see Appendix A.
@@
Table 1: Possible occupancy of the four degrees of freedom at $\nu=2$, yielding ${}_{4}C_{2}=6$ states. The charge Chern number $C_{c}$ sums the band Chern numbers, whereas the spin (valley) Chern number $C_{s}$ $(C_{v})$ represents the difference concerning the spin (valley). [Here, the band Chern number for the valley $K$ ($K^{\prime}$) is $+1$ ($-1$).] The three types of states – quantum anomalous Hall (QAH), quantum spin Hall (QSH), and quantum valley Hall (QVH) – are distinguished. The magnetization results in Fig. 3 are also summarized, with $\circ$ indicating finite magnetization. Note that our QSH state also supports the QVH effect [82].
@@ 6 $\times$ $\checkmark$ $\times$ $\checkmark$ (0,0,2)
Stacking graphene on hBN induces a moiré superlattice structure due to lattice mismatch. In this study, the twist angle is set to $0.77^{\circ}$ to align with the experimental conditions described in Ref. 5. Within the effective continuum model, the impact of hBN is represented by a local potential $v({\bm{r}})$ within the bottom graphene layer [60, 61] (see Appendix B for more details). The second quantized form of this potential is $V_{\text{hBN}}=\sum_{{\bm{k}}}\sum_{m_{1}m_{2}}\tilde{\bm{c}}^{\dagger}({\bm{k}}+m_{1}\bm{G}_{1}+m_{2}\bm{G}_{2})v(m_{1},m_{2})\tilde{\bm{c}}({\bm{k}})$, where $\tilde{\bm{c}}^{\dagger}=(c_{A_{1}}^{\dagger},c_{B_{1}}^{\dagger})$, $\bm{G}_{1}$ and $\bm{G}_{2}$ are the moiré reciprocal lattice vectors, and $v(m_{1},m_{2})$ is the Fourier coefficient of $v({\bm{r}})$. In the numerical calculations, the summation $\sum_{{\bm{k}}}\sum_{m_{1}m_{2}}$ is confined to the first and second moiré BZs.
@@
The effective noninteracting Hamiltonian is obtained as $H_{R5G}+V_{\text{hBN}}$. To incorporate many-body effects, the HF interaction is added. The original interaction term is $H_{\text{int}}=(1/2S)\sum_{{\bm{k}}{\bm{k}}^{\prime}{\bm{q}}}\sum_{ZZ^{\prime}}V_{C}({\bm{q}})c^{\dagger}_{Z}({\bm{k}}+{\bm{q}})c^{\dagger}_{Z^{\prime}}({\bm{k}}^{\prime}-{\bm{q}})c_{Z^{\prime}}({\bm{k}}^{\prime})c_{Z}({\bm{k}})$, where $S$ is the area of the system, and $Z$ and $Z^{\prime}$ represent the spin, valley, sublattice, and layer indices. The dual gate-screened Coulomb interaction used is $V_{C}({\bm{q}})=e^{2}/(2\epsilon_{0}\epsilon_{r}q)\tanh(qd_{s})$ for ${\bm{q}}\neq{\bm{0}}$ and $V_{C}({\bm{q}}={\bm{0}})=0$, with $d_{s}=25$ nm as the gate distance and $\epsilon_{r}=5$ as the dielectric constant [33]. More details on the HF interaction can be found in Appendix C. To reduce computational cost, the calculations are performed within the subspace of the first three noninteracting conduction bands for each valley and spin. The system with $24\times 24$ moiré unit cell is used unless otherwise specified.
III Band structure
@@
The self-consistent HF calculation assumes conservation of both spin and valley. The spin conservation assumption is valid due to the SU(2) spin-rotational symmetry of the HF Hamiltonian. Although valley conservation is nontrivial at this moment, we assume it based on the HF results at $\nu=1$ [35].
@@
Figure 2: Band structures at $\nu=2$ for the lowest energy Hartree-Fock solutions. (refer to Ref. [83]). Greek letters denote the high-symmetric points in the moiré BZ. Colors and line styles denote valley and spin properties. (a) and (b) correspond to the states #3 (QSH) and #1 (QAH) listed in Table 1, respectively. In the calculation for (b), a small “valley Zeeman term” is introduced to favor valley $K$ (refer to footnote [84]). The HF analysis is performed with the particle number fixed at $\nu=2$.
@@
Figures 2(a) and (b) display the HF band structures of lowest energy solutions, calculated by fixing the particle number at $\nu=2$. (To be precise, the state in Fig. 2(b) has a lower energy than that in (a) by ${}10^{-8}$eV per particle. Given the small difference, we consider the two states to be degenerate [83].) In Fig. 2(a), the occupied two lowest bands, distinctly separated from other conduction bands, exhibit $(K\!\!\uparrow,K^{\prime}\!\!\downarrow)$ with the band Chern numbers $C=(1,-1)$, corresponding to state #3 (QSH) in Table 1. In Fig. 2(b), the two lowest bands exhibit $(K\!\!\uparrow,K\!\!\downarrow)$ with band Chern numbers $C=(1,1)$, corresponding to state #1 (QAH) in Table 1. Because the HF Hamiltonian maintains symmetry for spin flipping in each valley, these two results indicate that the QAH, QSH, and QVH states in Table 1 are all degenerate ground states. [For example, state #5 (QVH) can be obtained by flipping the spin for $K^{\prime}$ in state #3.] These states do not emerge in the absence of the HF interaction. Their emergence results from many-body effects.
IV Lowest energy state